the sum of the length of the longer leg of a right triangle and 4 equals the length of the hypotenuse. The shorter leg is 4 inches shorter than the longer leg. Find the length of the longer leg.
Found 3 solutions by ikleyn, Alan3354, solver91311:
Answer by ikleyn(52780)

.
The legs are 12 and 16 inches long.

The hypotenuse is 20 inches long.

Classic (3,4,5)-right angled triangle.


Answer by Alan3354(69443)

the sum of the length of the longer leg of a right triangle and 4 equals the length of the hypotenuse. The shorter leg is 4 inches shorter than the longer leg. Find the length of the longer leg.
------------
a = longer leg, b = shorter. c = hypotenuse
----
b = a-4
c = a+4
----
(a-4)^2 + a^2 = (a+4)^2
2a^2 - 8a + 16 = a^2 + 8a + 16
a^2 - 16a = 0
a = 16




Answer by solver91311(24713)





Solve for


John

My calculator said it, I believe it, that settles it